OpenStudy (anonymous):

-5c/8=20 i think its -160, am i correct?

OpenStudy (anonymous):

Are you trying to find the value of C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok well -160 is incorrect its -32

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

Multiply both sides by 8. ((−5)(c)) 8 =20 ( ((−5)(c)) 8 )*(8)=(20)*(8) (−5)(c)=160 Step 2: Divide both sides by -5. −5c −5 = 160 −5 c=−32

OpenStudy (anonymous):

ok can i ask one more question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

-1/8 c = 2 i thought it was -16

OpenStudy (anonymous):

correct but positive

OpenStudy (anonymous):

why is it positive?

OpenStudy (anonymous):

wait -1/8 sorry

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

np